# Cholesky decomposition for symmetric matrices over finite fields

### v1

@@ Abstract.
Inspired by the seminal work of André-Louis Cholesky – whose contributions remain crucial even after more than a century in broader sciences – Cooper, Hanna and Whitlatch (2024) developed a theory of positive matrices over finite fields, and Khare and Vishwakarma (2025) described a general Cholesky factorization for a family of the dense cone of Hermitian matrices over real/complex fields, whose leading principal minors (LPM) are nonzero. Building on this, we develop a parallel theory within the finite field setting. Specifically (i) we extend the general Cholesky factorization to the LPM cone over finite fields which has asymptomatic density $1$. We show that (ii) this factorization is compatible with the entrywise Frobenius map, recently studied in the context of positivity preservers by Guillot, Gupta, Vishwakarma, and Yip [J. Algebra, 2025]. We also (iii) leverage the Cholesky-structures to define meaningful group operations on the matrix cone, and as an application (iv) enumerate sub-cones of LPM matrices using our general Cholesky factorizations.
@@ 1. Introduction and main results
For an integer $n\geqslant 1$, over a real or complex field $\mathbb{F}$, a Hermitian matrix $A\in\mathbb{F}^{n\times n}$ is called positive definite if the quadratic form $z^{*}Az>0$ for all nonzero $z\in\mathbb{F}^{n}$. Among more than half a dozen equivalent definitions of these matrices, André-Louis Cholesky’s Factorization Theorem stands out for its applicability in broader sciences and mathematics, even after one hundred years of its publication in 1924 [1]. While the applicability of Cholesky’s factorization of the positive definite cone of matrices has extensively been explored for over several decades now, there have been fewer attempts made to extend it to a broader class of matrices, and for fields that are not necessarily real or complex. We explore this extension over finite fields, and we begin by mentioning two such recent works [2, 7] that inspired the current article in such algebraic setting.
In a recent work by Cooper, Hanna, and Whitlatch [2], the cone of positive definite matrices over a finite field $\mathbb{F}_{q}$ (with $q$ elements) is introduced. The authors define a field element as positive if it is a nonzero square in the field. Then, a symmetric matrix with entries in $\mathbb{F}_{q}$ is called positive definite if all its leading principal minors are positive – that is, they are squares in the field. The authors show that such matrices admit a Cholesky-type factorization:
@@
While this result establishes a compelling analogue of Cholesky factorization in the finite field setting, other characterizations of positive definiteness familiar from the real or complex cases do not yet appear to have meaningful analogues over finite fields. Nonetheless, the existence of a Cholesky factorization in the finite field context is leveraged in [2] to demonstrate the existence of certain pressing sequences for weighted graphs. Furthermore, this work has served as a key motivation for developing a theory of positivity-preserving entrywise transformations over finite fields [4, 3], which we will discuss in more detail in a subsection later.
The second relevant work is more recent and is by Khare and Vishwakarma [7], who introduce a general Cholesky factorization and develop a rich theory for a dense set of real/complex Hermitian matrices. Their framework decomposes the space into cones of matrices determined by the sign patterns of their leading principal minors. Additionally, the framework explores several implications of this decomposition.
@@ 1. $(a)$
    We extend Theorem 1.1, which addresses positive definite matrices via Cholesky decomposition, to broader cones of symmetric matrices over a finite field.
@@ 2. $(b)$
    The general Cholesky decomposition discussed in [7] is done for fields (including the real-closed fields) in which the underlying/base field is totally ordered. In the current work over finite fields, the Cholesky decomposition is discussed over fields that do not have a total order, and hence demonstrating such factorizations in a different algebraic settings.
@@ 1.1. Notations
Each $LPM_{n}^{\mathbb{F}_{q}}(\epsilon)$ is nonempty, and so are $TPM_{n}^{\mathbb{F}_{q}}(\epsilon)$: define $\nabla,{\mathbb{D}}_{\epsilon}(\omega_{\pm})\in\mathbb{F}_{q}^{n\times n}$ via,
@@
The cones $LPM_{n}^{\mathbb{F}_{q}}$ and $TPM_{n}^{\mathbb{F}_{q}}$, defined as the set of all symmetric matrices with nonzero leading, and respectively trailing, principal minors, decompose into disjoint unions:
@@ 1.2. Cholesky decomposition and enumerations
where $\epsilon=\epsilon^{\prime}$ in the linear, and $\epsilon^{\prime}=(\epsilon_{n}\epsilon_{n-1},\epsilon_{n}\epsilon_{n-2},\dots,\epsilon_{n}\epsilon_{1},\epsilon_{n})$ in the nonlinear case.
@@
As an immediate corollary of Theorem A, we provide enumeration of the LPM and TPM (sub-)cones.
@@
Let $\mathrm{Sym}_{n}^{\mathbb{F}_{q}}$ denotes the set of all $n\times n$ symmetric matrices over a finite field $\mathbb{F}_{q}$. For a fixed $n\geqslant 1$, the proportion of matrices in $\mathrm{Sym}_{n}^{\mathbb{F}_{q}}$ that lie in the cone $LPM_{n}^{\mathbb{F}_{q}}$ (and similarly $TPM_{n}^{\mathbb{F}_{q}}$) tend to $1$ as $q\to\infty$; that is,
@@
It may be worth reiterating that the asymptotic density in question pertains to the setting where the matrix size $n$ is fixed, and the field size $q=p^{k}$ grows without bound, i.e., the integer $k\to\infty$.
@@ 1.3. The compatibility of Cholesky and Frobenius via Entrywise Transforms
Suppose $\mathbb{F}_{q}$ is a finite field, and consider a map $f:\mathbb{F}_{q}\to\mathbb{F}_{q}$. Then this map has a natural extension over all of the matrix space $\mathbb{F}_{q}^{n\times n}$, for integer $n\geqslant 1$ by considering:
@@
We now come to the possibilities over non-definite fields parallel to Theorem B. As we show later, the proof of Theorem B is an application of a key result from [3]. Other main results from [3] similarly yield a parallel statement for non-definite fields. (However, the compatibility question does not arise here as non-definite fields do not admit a Cholesky factorization.) We mention this result for completeness under a 33-byby-33 and a 22-byby-22 constraint. Note that slightly modified Remark 1.8(a) applies here.
@@
    The sign patterns are equal, and $f$ is a positive multiple of a power of Frobenius:
@@ 1.4. Two group structures over the LPM cones and their compatibility with the Frobenius map
Suppose we define two binary operations $\circledcirc_{1},\circledcirc_{2}:{\bf L}_{n}^{\mathbb{F}_{q}^{+}}\times{\bf L}_{n}^{\mathbb{F}_{q}^{+}}\to{\bf L}_{n}^{\mathbb{F}_{q}^{+}}$ by
@@
Remark 1.12 (Group structures on individual LPM cones and future directions).
@@
Furthermore, given a natural metric over a finite definite field, it would be interesting to investigate analogues of Riemannian geometric properties – similar to the setting explored in [7] over the real field.
@@
All results (and future directions) presented above in Subsections 1.3 and 1.4 for the LPM cones admit direct analogues for the TPM cones as well. Since the necessary modifications are straightforward, we omit the explicit statements in the interest of brevity.
@@ 2. $(2)$
    Collectively they show that for a given $A=\begin{pmatrix}B&{\bf b}\\ {\bf b}^{T}&c\end{pmatrix}\in LPM_{n}^{\mathbb{F}_{q}}(\epsilon)$, to obtain the required solution $L$ for $LA_{\epsilon}L^{T}=A$, we first need to be able to solve for a unique lower triangular $K\in\mathbb{F}_{q}^{(n-1)\times(n-1)}$ with positive diagonal entries, such that $B=KB_{\epsilon}K^{T}$. Assuming this can be solved, we show that ${\bf p}\in\mathbb{F}_{q}^{n-1}$ and $s\in\mathbb{F}_{q}^{+}$ can be solved uniquely from this. Note that if
@@
Now, the second half of the proof.
@@
The final steps of showing $(1)$ and $(2)$ for the TPM cones is a direct application of the assertions proved above for the linear and non-linear bijections. ∎
@@
The only step in the Proof of Theorem A(2) that does not work out for non-definite fields is the last paragraph of (2). More precisely, if $\mathbb{F}_{q}$ is non-definite, then for each $a\in\mathbb{F}_{q}^{+}$, there exists distinct $b_{1},b_{2}\in\mathbb{F}_{q}^{\pm}$ such that $b_{j}^{2}=a$. Therefore, the “positive square root” function here is not well-defined, and as a result a Cholesky factorization does not exist for non-definite fields.
@@
The implication $(2)\implies(1)$ follows from [3, Proposition 2.12]. For the other implication, suppose $A^{\prime}\in LPM_{s}^{\mathbb{F}_{q}}((1,1,\dots,1))$. Then $A:=A^{\prime}\oplus\mathbb{D}_{\epsilon^{\prime\prime}}\in LPM_{n}^{\mathbb{F}_{q}}(\epsilon^{\prime})$, where $\epsilon^{\prime\prime}=(\epsilon_{s+1},\dots,\epsilon_{n})$. Therefore $f[A]\in LPM_{n}^{\mathbb{F}_{q}}(\epsilon)$, and in particular, $f[A^{\prime}]=f[A]_{[s]\times[s]}\in LPM_{s}^{\mathbb{F}_{q}}((1,1,\dots,1))$. This means that f[]f[-] is an entrywise positivity preserver over $s\times s$ positive definite matrices, where $\mathbb{F}_{q}$ is a finite definite field. Therefore [3, Theorem B] implies that $f\equiv c\cdot\mathrm{Frob}^{\ell}$ for some $c\in\mathbb{F}_{q}^{+}$ and $\ell\in 0,1,\dots,k-1$. Moreover, it follows from [3, Proposition 2.12] that $\epsilon^{\prime}=\epsilon$. ∎
@@
Let $\mathbb{F}_{q}$ be a finite field with $q=p^{k}$ for $p$ a prime. Suppose $A$ and $B$ are matrices with entries in $\mathbb{F}_{q}$, such that $AB$ is well defined. Then $\mathrm{Frob}_{p}[AB]=\mathrm{Frob}_{p}[A]\mathrm{Frob}_{p}[A]$.

### v2

@@ Abstract.
Inspired by the seminal work of André-Louis Cholesky – whose contributions remain crucial in broader sciences even after more than a century – Cooper, Hanna and Whitlatch (2024) developed a theory of positive matrices over finite fields, and Khare and Vishwakarma (2025) described a general Cholesky factorization for a dense sub-family of the cone of Hermitian matrices over real/complex fields, whose leading principal minors (LPM) are nonzero. Building on this, we develop a parallel theory within the finite field setting. Specifically (i) we extend the general Cholesky factorization to the LPM cone over finite fields which has asymptotic density $1$. We show that (ii) this factorization is compatible with the entrywise Frobenius map, recently studied in the context of positivity preservers by Guillot, Gupta, Vishwakarma, and Yip [J. Algebra, 2025]. We also (iii) leverage the Cholesky-structures to define meaningful group operations on the matrix cone, and as an application (iv) enumerate sub-cones of LPM matrices using our general Cholesky factorizations.
@@ 1. Introduction and main results
For an integer $n\geqslant 1$, over the real or complex field $\mathbb{F}$, a Hermitian matrix $A\in\mathbb{F}^{n\times n}$ is called positive definite if the quadratic form $z^{*}Az>0$ for all nonzero $z\in\mathbb{F}^{n}$. Among more than half a dozen equivalent definitions of these matrices, the one given by André-Louis Cholesky’s Factorization Theorem stands out for its applicability in the broader sciences and mathematics, even after one hundred years of its publication in 1924 [1]. While the applicability of Cholesky’s factorization of the positive definite cone of matrices has extensively been explored for over several decades, there have been fewer attempts made to extend it to a broader class of matrices, and for fields that are not necessarily real or complex. We explore this extension over finite fields, and we begin by mentioning two such recent works [2, 7] that inspired the current article in such an algebraic setting.
In a recent work by Cooper, Hanna, and Whitlatch [2], the cone of positive definite matrices over a finite field $\mathbb{F}_{q}$ (with $q$ elements) is introduced. The authors define a field element to be positive if it is a nonzero square in the field. Then, a symmetric matrix with entries in $\mathbb{F}_{q}$ is called positive definite if all its leading principal minors are positive – that is, they are squares in the field. The authors show that such matrices admit a Cholesky-type factorization:
@@
While this result establishes a compelling analogue of the Cholesky factorization in the finite field setting, other characterizations of positive definiteness familiar from the real or complex cases do not yet appear to have meaningful analogues over finite fields. Nonetheless, the existence of a Cholesky factorization in the finite field context is leveraged in [2] to demonstrate the existence of certain pressing sequences for weighted graphs. Furthermore, this work has served as a key motivation for developing a theory of positivity-preserving entrywise transformations over finite fields [3, 4], which we will discuss in more detail in a subsection later.
The second relevant work is more recent and is by Khare and Vishwakarma [7], who introduce a general Cholesky factorization and develop a rich theory for a dense open set of real/complex Hermitian matrices. Their framework decomposes the space into cones of matrices determined by the sign patterns of their leading principal minors. Additionally, the framework explores several implications of this decomposition.
@@ 1. $(a)$
    We extend Theorem 1.1, which addresses positive definite matrices via their Cholesky decomposition, to broader cones of symmetric matrices over a finite field.
@@ 2. $(b)$
    The general Cholesky decomposition discussed in [7] is done for fields (including the real-closed fields) in which the underlying/base field is totally ordered. In the current work over finite fields, the Cholesky decomposition is discussed over fields that do not have a total order, thus demonstrating such factorizations in a different algebraic setting.
@@ 1.1. Notations
Each $LPM_{n}^{\mathbb{F}_{q}}(\epsilon)$ is nonempty, and so is $TPM_{n}^{\mathbb{F}_{q}}(\epsilon)$: define $\nabla,{\mathbb{D}}_{\epsilon}(\omega_{\pm})\in\mathbb{F}_{q}^{n\times n}$ via,
@@
The cones $LPM_{n}^{\mathbb{F}_{q}}$ and $TPM_{n}^{\mathbb{F}_{q}}$, defined as the sets of all symmetric matrices with nonzero leading, and respectively trailing, principal minors, decompose into disjoint unions:
@@ 1.2. Cholesky decomposition and enumerations
where $\epsilon=\epsilon^{\prime}$ in the linear case, and $\epsilon^{\prime}=(\epsilon_{n}\epsilon_{n-1},\epsilon_{n}\epsilon_{n-2},\dots,\epsilon_{n}\epsilon_{1},\epsilon_{n})$ in the nonlinear case.
@@
As an immediate corollary of Theorem A, we provide enumerations of the LPM and TPM (sub-)cones.
@@
Let $\mathrm{Sym}_{n}^{\mathbb{F}_{q}}$ denotes the set of all $n\times n$ symmetric matrices over a finite field $\mathbb{F}_{q}$. For a fixed $n\geqslant 1$, the proportion of matrices in $\mathrm{Sym}_{n}^{\mathbb{F}_{q}}$ that lie in the cone $LPM_{n}^{\mathbb{F}_{q}}$ (and similarly $TPM_{n}^{\mathbb{F}_{q}}$) tends to $1$ as $q\to\infty$; that is,
@@
It may be worth reiterating that the asymptotic density in question pertains to the setting where the matrix size $n$ is fixed, and the field size $q=p^{k}$ grows without bound (e.g., the integer $k\to\infty$ if we fix $p$).
@@ 1.3. The compatibility of Cholesky and Frobenius via Entrywise Transforms
Suppose $\mathbb{F}_{q}$ is a finite field, and consider a map $f:\mathbb{F}_{q}\to\mathbb{F}_{q}$. Then this map has a natural extension to all of the matrix space $\mathbb{F}_{q}^{n\times n}$, for any integer $n\geqslant 1$ by considering:
@@
We now come to the possibilities over non-definite fields parallel to Theorem B. As we show later, the proof of Theorem B is an application of a key result from [3]. Other main results from [3] similarly yield a parallel statement for non-definite fields. (However, the compatibility question does not arise here as non-definite fields do not admit a Cholesky factorization.) We mention this result for completeness under a 33-byby-33 and a 22-byby-22 constraint. Note that a slightly modified Remark 1.8(a) applies here.
@@ 1.4. Two group structures over the LPM cones and their compatibility with the Frobenius map
Define two binary operations $\circledcirc_{1},\circledcirc_{2}:{\bf L}_{n}^{\mathbb{F}_{q}^{+}}\times{\bf L}_{n}^{\mathbb{F}_{q}^{+}}\to{\bf L}_{n}^{\mathbb{F}_{q}^{+}}$ by
@@
Remark 1.12 (Group structures on individual LPM cones).
@@
All results (and future directions) presented above in Subsections 1.3 and 1.4 for the LPM cones admit direct analogues for the TPM cones as well. Since the necessary modifications are straightforward, we omit the explicit statements and proofs in the interest of brevity.
@@ 2. $(2)$
    Collectively, they show that for a given $A=\begin{pmatrix}B&{\bf b}\\ {\bf b}^{T}&c\end{pmatrix}\in LPM_{n}^{\mathbb{F}_{q}}(\epsilon)$, to obtain the required solution $L$ for $LA_{\epsilon}L^{T}=A$, we first need to be able to solve for a unique lower triangular $K\in\mathbb{F}_{q}^{(n-1)\times(n-1)}$ with positive diagonal entries, such that $B=KB_{\epsilon}K^{T}$. Assuming this can be solved, we show that ${\bf p}\in\mathbb{F}_{q}^{n-1}$ and $s\in\mathbb{F}_{q}^{+}$ can be solved uniquely from this. Note that if
@@
Now for the the second half of the proof.
@@
The final steps of showing $(1)$ and $(2)$ for the TPM cones are a direct applications of the assertions proved above for the linear and non-linear bijections. ∎
@@
The only step in the Proof of Theorem A(2) that does not work out for non-definite fields is the last paragraph of (2). More precisely, if $\mathbb{F}_{q}$ is non-definite, then for each $a\in\mathbb{F}_{q}^{+}$, there exist distinct $b_{1},b_{2}\in\mathbb{F}_{q}^{\pm}$ such that $b_{j}^{2}=a$. Therefore, the “positive square root” function is not well-defined here, and as a result a Cholesky factorization does not exist for non-definite fields.
@@
The implication $(2)\implies(1)$ follows from [3, Proposition 2.12]. For the other implication, suppose $A^{\prime}\in LPM_{s}^{\mathbb{F}_{q}}((1,1,\dots,1))$. Then $A:=A^{\prime}\oplus\mathbb{D}_{\epsilon^{\prime\prime}}(\omega_{\pm})\in LPM_{n}^{\mathbb{F}_{q}}(\epsilon^{\prime})$, for some $\omega_{\pm}\in\mathbb{F}_{q}^{\pm}$, where $\epsilon^{\prime\prime}=(\epsilon_{s+1},\dots,\epsilon_{n})$. Therefore $f[A]\in LPM_{n}^{\mathbb{F}_{q}}(\epsilon)$, and in particular, $f[A^{\prime}]=f[A]_{[s]\times[s]}\in LPM_{s}^{\mathbb{F}_{q}}((1,1,\dots,1))$. This means that f[]f[-] is an entrywise positivity preserver over $s\times s$ positive definite matrices, where $\mathbb{F}_{q}$ is a finite definite field. Therefore [3, Theorem B] implies that $f\equiv c\cdot\mathrm{Frob}^{\ell}$ for some $c\in\mathbb{F}_{q}^{+}$ and $\ell\in 0,1,\dots,k-1$. Moreover, it follows from [3, Proposition 2.12] that $\epsilon^{\prime}=\epsilon$. ∎
@@
Let $\mathbb{F}_{q}$ be a finite field with $q=p^{k}$ for $p$ a prime. Suppose $A$ and $B$ are matrices with entries in $\mathbb{F}_{q}$, such that $AB$ is well defined. Then $\mathrm{Frob}_{p}[AB]=\mathrm{Frob}_{p}[A]\mathrm{Frob}_{p}[B]$.